# Principal stratification with $U$-statistics under principal ignorability

### v2

@@ Abstract
Principal stratification is a popular framework for causal inference in the presence of an intermediate outcome. While the principal average treatment effects have traditionally been the default target of inference, it may not be sufficient when the interest lies in the relative favorability of one potential outcome over the other within the principal stratum. We thus introduce the principal generalized causal effect estimands, which extend the principal average causal effects to accommodate nonlinear contrast functions. Under principal ignorability, we expand the theoretical results in Jiang et al., (2022) to a much wider class of causal estimands in the presence of a binary intermediate variable. We develop identification formulas and derive the efficient influence functions of the generalized estimands for principal stratification analyses. These efficient influence functions motivate a set of multiply robust estimators and lay the ground for obtaining efficient debiased machine learning estimators via cross-fitting based on $U$-statistics. The proposed methods are illustrated through simulations and the analysis of a data example.
Keywords: causal inference, efficient influence function, principal stratification, multiply robust estimation, $U$-statistics, win-ratio.
@@ 1 Introduction
For causal inference under the Neyman-Rubin model (Rubin,, 1974), the average causal effect—defined as the expected contrast between two potential outcomes under alternative treatment conditions—is often the primary estimand of interest. Specifically, let $\{Y(1),Y(0)\}$ denote a pair of potential outcomes for a randomly selected individual from a super-population, and then the average causal effect estimand is defined as $\mathbb{E}\{Y(1)-Y(0)\}$. Conceptually, the average causal effect estimand can be expanded to a wider class, allowing arbitrary comparisons. That is, given a contrast function $h(u,v)$ defined on the product space, $\mathcal{Y}\times\mathcal{Y}$ ($\mathcal{Y}$: the potential outcome space), a generalized estimand can be defined as $\mathbb{E}[h\{Y(1),Y(0)\}]$. Examples include the average causal effect when $h(u,v)=u-v$, and the proportion of the population benefiting from treatment, $\mathbb{P}\{Y(1)>Y(0)\}$, when $h(u,v)=\mathbbm{1}(u>v)$ (Lu et al.,, 2020). Because of the fundamental problem of causal inference, $\mathbb{E}[h\{Y(1),Y(0)\}]$ is generally inestimable for nonlinear $h$. However, a closely related quantity, the generalized causal effect (O’brien,, 1988; Mao,, 2017), can be formulated as:
@@
where $F_{z}(\cdot)$ is the marginal distribution of the potential outcome $Y(z)\in\mathcal{Y}$ for $z=0,1$. Consider $\{Y_{1}(1),Y_{1}(0)\}$ and $\{Y_{2}(1),Y_{2}(0)\}$ as two independent replicates of $\{Y(1),Y(0)\}$, and then $\tau_{h}$ can be re-expressed in a familiar form, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}]$. Two notable examples are the probability index (the target parameter for the Mann-Whitney statistic), $\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)\}$, when $h(u,v)=\mathbbm{1}(u\geq v)$ (Vermeulen et al.,, 2015), and the win-ratio, $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)\}/\mathbb{P}\{Y_{1}(1)<Y_{2}(0)\}$, when $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (Pocock et al.,, 2011). These estimable quantities are alternative measurements for the population causal effect and can be straightforwardly interpreted. For example, the probability index comparing the ranking of outcomes under two potential outcome distributions measures the probability that a randomly chosen unit from $F_{1}$ has a higher-ranked outcome than $F_{0}$, and is robust against extreme values in the tails of the potential outcome distributions.
@@
When drawing causal inferences from randomized experiments or observational studies, intermediate variables such as treatment noncompliance (Angrist et al.,, 1996), mortality event (Zhang et al.,, 2009) or dropout (Frangakis and Rubin,, 1999), often have important implications even on the definition of a causal effect and require careful adjustment. The intermediate variable is not guaranteed to be balanced via randomization, and direct conditioning on the intermediate variable leads to sample selection bias and non-causal comparisons even in randomized experiments. As a solution, the principal stratification framework (Frangakis and Rubin,, 2002) segments the population into subpopulations (principal strata) according to joint potential values of the intermediate variable under alternative conditions. The stratum-specific causal effect—the principal causal effect—can be defined and is causally interpretable. However, a predominant focus in prior works has been on the average causal effect within each principal stratum, and there is a lack of related development when the generalized causal effect is of interest within each subpopulation. Therefore, in this article, we propose a new class of estimands, the principal generalized causal effect (PGCE), which accommodates an arbitrary and potentially nonlinear contrast function. Under a linear contrast function, the PGCE reduces to the usual principal average causal effect. As we demonstrate in due course, the PGCE can represent the stratum-specific probability index and strata-specific win-ratio with the appropriate choice of the contrast function, and it can provide a meaningful assessment of the relative treatment effect within important subpopulations.
@@
Assuming a binary intermediate variable, we study the identification and inference of the PGCE under two structural assumptions, monotonicity (Angrist et al.,, 1996) and principal ignorability (Jo and Stuart,, 2009; Ding and Lu,, 2016), which are typically invoked for principal stratification analyses. The monotonicity assumption states that the treatment has a nonnegative causal effect on the intermediate outcome of each unit, effectively restricting the number of potential strata. This condition is considered plausible when, for example, the intermediate variable is death, and the treatment is known to benefit survival or when the intermediate variable is the actual treatment received and the treatment does not discourage treatment uptake. The principal ignorability assumption, on the other hand, requires that the observed pre-treatment covariates are sufficient to control for confounding between the intermediate variable and the final outcome. For the principal average causal effect, under monotonicity and principal ignorability, Ding and Lu, (2016) formalized weighting estimators using principal scores (the conditional probabilities of the latent principal strata given observed covariates); Jiang et al., (2022) provided alternative identification formulas based on the treatment propensity score, the principal score and the potential outcome models, and further developed multiply robust and locally efficient estimators under the additional assumption of treatment unconfoundedness. In contrast, the identification and the semiparametric efficiency theory for the PGCE in the presence of a binary intermediate variable remain largely unclear.
@@
The contributions of this article are several-folded. First, we propose a class of PGCE estimands in the presence of an intermediate variable, and develop complementary nonparametric identification formulas under treatment unconfoundedness, monotonicity, and principal ignorability. Second, we provide a formal treatment of semiparametric efficiency theory for estimating the PGCE estimands, and seek efficiency improvement by deriving their corresponding efficient influence functions (Bickel et al.,, 1993, EIFs). Specifically, we show that, when the contrast function $h(u,v)=u-v$, our EIFs reduce to those derived by Jiang et al., (2022) as a special case, but are otherwise more general to accommodate arbitrary contrast functions. A major complexity in the context of a nonlinear contrast function, however, is that the estimating equations based on the EIFs requires nuisance functions defined over pairs of observations. Thus, to obtain a feasible estimator, we define ratio $U$-statistics whose influence functions are the same as the EIFs. We show that the $U$-estimators are triply robust in the sense that they are consistent and asymptotically normal when two of the three nuisance functions (including the treatment propensity score, principal score, and the potential outcome models) are correctly specified. When all nuisance functions are correctly specified, the $U$-estimators are locally efficient, with their asymptotic variances achieving the semiparametric efficiency lower bounds. The asymptotic properties of the $U$-estimators are developed upon the theory of empirical process (van der Vaart,, 1998, §19), $U$-process (Arcones and Gine,, 1993), as well as the theory of $U$-statistics (van der Vaart,, 1998, §11-12). Finally, the construction of the multiply robust $U$-estimators facilitates the integration of machine learning estimators for the nuisance functions. To achieve full asymptotic efficiency, we learn each nuisance function using machine learning methods, and operationalize the debiased machine learning $U$-estimators via cross-fitting (Chernozhukov et al.,, 2018). We discuss the modification needed for cross-fitting when a nonlinear contrast function is considered when defining the PGCE. We prove the multiple robustness and efficiency property of the debiased machine learning estimators under mild regulation conditions (Chernozhukov et al.,, 2022; Escanciano and Terschuur,, 2023).
@@
The rest of this article is organized as follows. Section 2 introduces our causal estimands and identification assumptions for principal stratification. Section 3 discusses the nonparametric identification formulas for the PGCE and associated moment estimators. Section 4 derives the EIF for each PGCE estimand, and develops the EIF-motivated feasible $U$-estimators and the debiased machine learning estimators. Section 5 demonstrates the finite-sample performance of the proposed estimators via simulations. Section 6 illustrates the proposed methods with a data example from the United States Jobs Corps study (Schochet et al.,, 2001), and Section 7 concludes. Detailed proofs of all Theorems are available in the Online Supplement.
@@ 2.1 Setting and notation
We consider a study (experimental or observational) with $n$ participants and two treatment conditions, and let $Z_{i}$ denote the treatment assignment of participant $i\in\{1,\ldots,n\}$; $Z_{i}=1$ indicates the treatment condition and 00 the control. We use $D_{i}(z)$ and $Y_{i}(z)$ to denote the potential outcomes of the intermediate variable and the final outcome variable of interest for participant $i$ under treatment $Z_{i}=z$ ($z\in\{0,1\}$), respectively. The intermediate variable $D_{i}\in\{0,1\}$ is assumed to be binary, but there is no restriction on the type of variable for $Y_{i}$. Under the stable unit treatment value assumption (Rubin,, 1980, SUTVA), the observed intermediate and final outcomes can be represented as $D_{i}=Z_{i}D_{i}(Z_{i})+(1-Z_{i})D(1-Z_{i})$ and $Y_{i}=Z_{i}Y_{i}(Z_{i})+(1-Z_{i})Y(1-Z_{i})$. Let $X_{i}$ represent the vector of observed pre-treatment covariates, and the observed data for participant $i$ is thus denoted by $\mathcal{O}_{i}=\{X_{i},Z_{i},D_{i},Y_{i}\}$, where $\mathcal{O}_{i}$’s are assumed to be independent and identically distributed (i.i.d).
@@ 2.2 Principal generalized causal effects
We operate under the principal stratification framework (Frangakis and Rubin,, 2002), and define distinct principal strata according to the joint potential values of the intermediate variable under alternative conditions. Specifically, we use $S=\{D(1),D(0)\}$ to denote the principal stratum variable, which has four potential values: $\{11,10,01,00\}$. A key insight under the principal stratification framework is that, although $D$ is a post-randomization variable, the strata variable $S$ is unaffected by the assignment and can be viewed as a pre-treatment variable. Therefore, based on any contrast function $h$, we define the principal generalized causal effect estimand as
@@
where $Y_{1}(1)$ and $Y_{2}(0)$ denote the potential outcomes for two randomly selected units from the subpopulation with strata membership $s$ ($S_{1}=s$ and $S_{2}=s$ denote the strata membership for these two selected units). Of note, to ensure a proper causal interpretation of $\tau_{h}^{s}$, it is required that both elements of the pair for comparison are from the same principal stratum such that the comparison is made within a common subpopulation. If the two units are selected from different principal strata, the resulting quantity, $\tau_{h}^{s_{1},s_{2}}=\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=s_{1},S_{2}=s_{2}]$ with $s_{1}\neq s_{2}$, remains estimable under this framework but generally lacks a meaningful causal interpretation; we thus do not further pursue this direction. Evidently, when a linear contrast function is specified as $h(u,v)=u-v$, the PGCE in (2) reduces to $\tau_{h}^{s}=\mathbb{E}\{Y(1)-Y(0)|S=s\}$, which is precisely the principal average causal effect studied in the existing principal stratification literature; see, for example, Frangakis and Rubin, (2002), Jiang et al., (2022) and Lipkovich et al., (2022). When a nonlinear contrast function is considered in defining the PGCE, $\tau_{h}^{s}$ is no longer a simple average of $Y(z)$ within each principal stratum. The following two examples illustrate the PGCE with a specific contrast function.
Example 2.1 (Probability index under truncation by death).
When $D$ represents the survival status (alive if $D=1$) that is ascertained before the measurement of the final outcome $Y$ (such as quality of life, which is only well-defined when $D=1$), there exist four principal strata based on the potential survival status. That is, $s=11$ represents the always-survivors who would survive until the measurement of the final outcome regardless of the assignment; $s=10$ and $s=01$ represent the protected and harmed, who would survive until the measurement of the final outcome only under the treatment condition and control condition, respectively; $s=00$ represents the never-survivors who would not survive until the measurement of the final outcome regardless of the assignment. Because the final outcome is oftentimes only well-defined without ambiguity when $D=1$, the survivor average causal effect (SACE) has been suggested as a relevant target of inference with $\mathrm{SACE}=\mathbb{E}\{Y(1)-Y(0)|S=11\}$ (Zhang et al.,, 2009; Chen et al.,, 2024). In clinical research, Acion et al., (2006) argued that a more intuitive effect size can be measured by the probability index defined with a contrast function $h(u,v)=\mathbbm{1}(u\geq v)$. Based on this contrast function, we can define the survivor probability index (SPI) as
@@
Compared to the SACE, the SPI is less sensitive to skewness of the underlying distributions of the potential outcomes and heteroskedasticity between conditions and outliers. In the special case of normality and under the equal variance assumption, the SPI is a strictly monotone function of the SACE because $\mathrm{SPI}=\Phi(\mathrm{SACE}/\sqrt{2}\sigma)$, where $\Phi$ is the standard Gaussian cumulative distribution function and $\sigma^{2}$ is the constant variance of the potential outcomes (Wolfe and Hogg,, 1971).
Example 2.2 (Win estimands under noncompliance).
In experimental studies with ordinal outcomes, Wang and Pocock, (2016) introduced the win-ratio estimand for pairwise comparisons, $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)\}/\allowbreak\mathbb{P}\{Y_{1}(1)<Y_{2}(0)\}$. This estimand corresponds to a two-dimensional contrast function specified as $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$. If $D$ represents the actual treatment received that can differ from the assignment, then $S\in\{11,10,01,00\}$ refers to always-takers, compliers, defiers, and never-takers, respectively (Angrist et al.,, 1996). Therefore, the win-ratio estimands can be extended to define the compliers win-ratio (CWR) as
@@
The CWR estimand represents the odds of a randomly selected unit from all-treated compliers having a more favorable outcome than one from the all-control compliers subpopulation. Alternatively, the complier win-difference (CWD) or complier net benefit estimand on the difference scale is given as
@@
The win estimands within other principal strata can be defined similarly. Mao, (2024) considered the point and interval identification of a similar estimand using the instrumental variable approach.

### v3

@@ Abstract
Principal stratification is a popular framework for causal inference in the presence of an intermediate outcome. While the principal average treatment effects have traditionally been the default target of inference, it may not be sufficient when the interest lies in the relative favorability of one potential outcome over the other within the principal stratum. We thus introduce the principal generalized causal effect estimands, which extend the principal average causal effects to accommodate nonlinear contrast functions. Under principal ignorability, we expand the theoretical results in Jiang et al. (2022) to a much wider class of causal estimands in the presence of a binary intermediate variable. We develop identification formulas and derive the efficient influence functions of the generalized estimands for principal stratification analyses. These efficient influence functions motivate a set of multiply robust estimators and lay the ground for obtaining efficient debiased machine learning estimators via cross-fitting based on $U$-statistics. The proposed methods are illustrated through simulations and the analysis of a data example.
Keywords: Causal inference, efficient influence function, principal stratification, multiply robust estimation, $U$-statistics, win ratio
@@ 1 Introduction
For causal inference under the Neyman-Rubin model (Rubin, 1974), the average causal effect—defined as the expected contrast between two potential outcomes under alternative treatment conditions—is often the primary estimand of interest. Specifically, let $\{Y(1),Y(0)\}$ denote a pair of potential outcomes for a randomly selected individual from a super-population, and then the average causal effect estimand is defined as $\mathbb{E}\{Y(1)-Y(0)\}$. Conceptually, the average causal effect estimand can be expanded to a wider class, allowing arbitrary comparisons. That is, given a contrast function $h(u,v)$ defined on the product space, $\mathcal{Y}\times\mathcal{Y}$ ($\mathcal{Y}$: the potential outcome space), a generalized estimand can be defined as $\mathbb{E}[h\{Y(1),Y(0)\}]$. Examples include the average causal effect when $h(u,v)=u-v$, and the proportion of the population benefiting from treatment, $\mathbb{P}\{Y(1)>Y(0)\}$, when $h(u,v)=\mathbbm{1}(u>v)$ (Lu et al., 2020). Because $Y(1)$ and $Y(0)$ are never observed simultaneously, $\mathbb{E}[h\{Y(1),Y(0)\}]$ is generally inestimable for nonlinear $h$. However, a closely related quantity, the generalized causal effect (O’brien, 1988; Mao, 2017), can be formulated as:
@@
where $F_{z}(\cdot)$ is the marginal distribution of the potential outcome $Y(z)\in\mathcal{Y}$ for $z=0,1$. Let $\{Y_{1}(1),Y_{1}(0)\}$ and $\{Y_{2}(1),Y_{2}(0)\}$ denote two independent replicates of $\{Y(1),Y(0)\}$, and $\tau_{h}$ can be re-expressed in a familiar form, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}]$. Two notable examples are the probability index (the target parameter for the Mann-Whitney statistic), $\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)\}$, when $h(u,v)=\mathbbm{1}(u\geq v)$ (Vermeulen et al., 2015), and the win ratio, $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)\}/\mathbb{P}\{Y_{1}(1)<Y_{2}(0)\}$, when $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (Pocock et al., 2011). These estimable quantities are alternative measurements for the population causal effect and can be straightforwardly interpreted. For example, the probability index comparing the ranking of outcomes under two potential outcome distributions measures the probability that a randomly chosen unit from $F_{1}$ has a higher-ranked outcome than $F_{0}$ and is robust against extreme values in the tails of the potential outcome distributions.
@@
When drawing causal inferences from randomized experiments or observational studies, intermediate variables such as treatment noncompliance (Angrist et al., 1996), mortality event (Zhang et al., 2009) or dropout (Frangakis and Rubin, 1999), often have important implications even on the definition of a causal effect and require careful adjustment. The intermediate variable is not guaranteed to be balanced via randomization, and direct conditioning on the intermediate variable leads to sample selection bias and non-causal comparisons, even in randomized experiments. As a solution, the principal stratification framework (Frangakis and Rubin, 2002) segments the population into principal strata (subpopulations) according to joint potential values of the intermediate variable under alternative conditions. The stratum-specific causal effect—the principal causal effect—can thus be defined and is causally interpretable. However, a predominant focus in prior works has been on the average causal effect within each principal stratum, and there is a lack of related development when the generalized causal effect is of interest within each subpopulation. Therefore, in this article, we propose a new class of estimands, the principal generalized causal effect (PGCE), which accommodates an arbitrary and potentially nonlinear contrast function. The PGCE, under a linear contrast function, reduces to the usual principal average causal effect. As we demonstrate in due course, the PGCE can represent the stratum-specific probability index and strata-specific win ratio with the appropriate choice of the contrast function, and it can provide a meaningful assessment of the relative treatment effect within subpopulations of interest.
@@
Assuming a binary intermediate variable, we study the identification and inference of the PGCE under two structural assumptions, monotonicity (Angrist et al., 1996) and principal ignorability (Jo and Stuart, 2009; Ding and Lu, 2016), which are typically invoked for principal stratification analyses. The monotonicity assumption states that the treatment has a nonnegative causal effect on the intermediate outcome of each unit, effectively restricting the number of potential strata. This condition is considered plausible when, for example, the intermediate variable is death, and the treatment is known to benefit survival, or when the intermediate variable is the actual treatment received and the treatment does not discourage treatment uptake. The principal ignorability assumption, on the other hand, requires that the observed pre-treatment covariates are sufficient to control for confounding between the intermediate variable and the final outcome. For the principal average causal effect, under monotonicity and principal ignorability, Ding and Lu (2016) formalized weighting estimators using principal scores (the conditional probabilities of the latent principal strata given observed covariates); Jiang et al. (2022) provided alternative identification formulas based on the treatment propensity score, the principal score and the potential outcome models, and further developed multiply robust and locally efficient estimators under the additional assumption of treatment unconfoundedness. In contrast, the identification and the semiparametric efficiency theory for the PGCE in the presence of a binary intermediate variable remain largely unclear.
@@
The contributions of this article are several-fold. First, we propose a class of PGCE estimands in the presence of an intermediate variable and develop complementary nonparametric identification formulas under treatment unconfoundedness, monotonicity, and principal ignorability. Second, we provide a formal treatment of semiparametric efficiency theory for estimating the PGCE estimands and seek efficiency improvement by deriving their corresponding efficient influence functions (Bickel et al., 1993, EIFs). Specifically, we show that when the contrast function $h(u,v)=u-v$, our EIFs reduce to those derived by Jiang et al. (2022) as a special case but are otherwise more general to accommodate arbitrary contrast functions. A major complexity in the context of a nonlinear contrast function, however, is that the estimating equations based on the EIFs require nuisance functions defined over pairs of observations. Thus, to obtain a feasible estimator, we define ratio $U$-statistics whose influence functions are the same as the EIFs. We show that the $U$-estimators are triply robust in the sense that they are consistent and asymptotically normal when two of the three nuisance functions (including the treatment propensity score, principal score, and the potential outcome models) are correctly specified. When all nuisance functions are correctly specified, the $U$-estimators are locally efficient, with their asymptotic variances achieving the semiparametric efficiency lower bounds. The asymptotic properties of the $U$-estimators are developed upon the theory of empirical process (van der Vaart, 1998, §19), $U$-process (Arcones and Gine, 1993), as well as the theory of $U$-statistics (van der Vaart, 1998, §11-12). Finally, the construction of the multiply robust $U$-estimators facilitates the integration of machine learning estimators for the nuisance functions. To achieve full asymptotic efficiency, we learn each nuisance function using machine learning methods and operationalize the debiased machine learning $U$-estimators via cross-fitting (Chernozhukov et al., 2018). We discuss the modification needed for cross-fitting if a nonlinear contrast function is considered when defining the PGCE and prove the multiple robustness and efficiency property of the debiased machine learning estimators under mild regulation conditions (Chernozhukov et al., 2022; Escanciano and Terschuur, 2023).
@@
The rest of this article is organized as follows. Section 2 introduces our causal estimands and identification assumptions for principal stratification. Section 3 discusses the nonparametric identification formulas for the PGCE and associated moment estimators. Section 4 derives the EIF for each PGCE estimand and develops the EIF-motivated feasible $U$-estimators and the debiased machine learning estimators. Section 5 demonstrates the finite-sample performance of the proposed estimators via simulations. Section 6 illustrates the proposed methods with a data example from the United States Jobs Corps study (Schochet et al., 2001), and Section 7 concludes. Detailed proofs of all theoretical results are available in the Online Supplement.
@@ 2.1 Setting and notation
We consider a study (experimental or observational) with $n$ participants and two treatment conditions and let $Z_{i}$ denote the treatment assignment of participant $i\in\{1,\ldots,n\}$; $Z_{i}=1$ indicates the treatment condition and 00 the control. We use $D_{i}(z)$ and $Y_{i}(z)$ to denote the potential outcomes of the intermediate variable and the final outcome variable of interest for participant $i$ under treatment $Z_{i}=z$ ($z\in\{0,1\}$), respectively. The intermediate variable $D_{i}\in\{0,1\}$ is assumed to be binary, but there is no restriction on the type of variable for $Y_{i}$. Under the stable unit treatment value assumption (Rubin, 1980, SUTVA), the observed intermediate and final outcomes can be represented as $D_{i}=Z_{i}D_{i}(Z_{i})+(1-Z_{i})D(1-Z_{i})$ and $Y_{i}=Z_{i}Y_{i}(Z_{i})+(1-Z_{i})Y(1-Z_{i})$. Let $X_{i}$ represent the vector of observed pre-treatment covariates, and the observed data for participant $i$ is thus denoted by $\mathcal{O}_{i}=\{X_{i},Z_{i},D_{i},Y_{i}\}$, where $\mathcal{O}_{i}$’s are assumed to be independent and identically distributed (i.i.d).
@@ 2.2 Principal generalized causal effects
We operate under the principal stratification framework (Frangakis and Rubin, 2002), and define distinct principal strata according to the joint potential values of the intermediate variable under alternative conditions. Specifically, we use $S=\{D(1),D(0)\}$ to denote the principal stratum variable, which has four potential values: $\{11,10,01,00\}$. A key insight under the principal stratification framework is that although $D$ is a post-randomization variable, the stratum variable $S$ is unaffected by the assignment and can be viewed as a pre-treatment variable. Therefore, based on any contrast function $h$, we define the principal generalized causal effect (PGCE) estimand as
@@
where $Y_{1}(1)$ and $Y_{2}(0)$ denote the potential outcomes for two randomly selected units from the subpopulation with strata membership $s$ ($S_{1}=s$ and $S_{2}=s$ denote the strata membership for these two selected units). Of note, to ensure a proper causal interpretation of $\tau_{h}^{s}$, it is required that both elements of the pair for comparison are from the same principal stratum such that the comparison is made within a common subpopulation. If the two units are selected from different principal strata, the resulting quantity, $\tau_{h}^{s_{1},s_{2}}=\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=s_{1},S_{2}=s_{2}]$ with $s_{1}\neq s_{2}$, remains estimable under this framework but generally lacks a meaningful causal interpretation; we thus do not further pursue this direction. Evidently, when a linear contrast function is specified as $h(u,v)=u-v$, the PGCE in (2) reduces to $\tau_{h}^{s}=\mathbb{E}\{Y(1)-Y(0)|S=s\}$, which is precisely the principal average causal effect studied in the existing principal stratification literature; see, for example, Frangakis and Rubin (2002), Jiang et al. (2022) and Lipkovich et al. (2022). When a nonlinear contrast function is considered in defining the PGCE, $\tau_{h}^{s}$ is no longer a simple average of $Y(z)$ within each principal stratum. The following two examples illustrate the PGCE with a specific contrast function.
Example 1 (Probability index under truncation by death)
When $D$ represents the survival status (alive if $D=1$) that is ascertained before the measurement of the final outcome $Y$ (such as quality of life, which is only well-defined when $D=1$), there exist four principal strata based on the potential survival status. That is, $s=11$ represents the always-survivors who would survive until the measurement of the final outcome regardless of the assignment; $s=10$ and $s=01$ represent the protected and harmed, who would survive until the measurement of the final outcome only under the treatment condition and control condition, respectively; $s=00$ represents the never-survivors who would not survive until the measurement of the final outcome regardless of the assignment. Because the final outcome is oftentimes only well-defined without ambiguity when $D=1$, the survivor average causal effect (SACE) has been suggested as a relevant target of inference with $\mathrm{SACE}=\mathbb{E}\{Y(1)-Y(0)|S=11\}$ (Zhang et al., 2009; Chen et al., 2024). In clinical research, Acion et al. (2006) argued that a more intuitive effect size can be measured by the probability index defined with a contrast function $h(u,v)=\mathbbm{1}(u\geq v)$. Based on this contrast function, we can define the survivor probability index (SPI) as
@@
Compared to the SACE, the SPI is less sensitive to the skewness of the underlying distributions of the potential outcomes and heteroskedasticity between conditions and outliers. In the special case of normality and under the equal variance assumption, the SPI is a strictly monotone function of the SACE because $\mathrm{SPI}=\Phi(\mathrm{SACE}/\sqrt{2}\sigma)$, where $\Phi$ is the standard Gaussian cumulative distribution function and $\sigma^{2}$ is the constant variance of the potential outcomes (Wolfe and Hogg, 1971).
Example 2 (Win estimands under noncompliance)
In experimental studies with ordinal outcomes, Wang and Pocock (2016) introduced the win ratio estimand for pairwise comparisons, $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)\}/\mathbb{P}\{Y_{1}(1)<Y_{2}(0)\}$. This estimand corresponds to a two-dimensional contrast function specified as $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$. If $D$ represents the actual treatment received that can differ from the assignment, then $S\in\{11,10,01,00\}$ refers to always-takers, compliers, defiers, and never-takers, respectively (Angrist et al., 1996). Therefore, the win ratio estimands can be extended to define the compliers win ratio (CWR) as
@@
The CWR estimand represents the odds of a randomly selected unit from all-treated compliers having a more favorable outcome than one from the all-control compliers subpopulation. Alternatively, the complier win difference (CWD) or complier net benefit estimand on the difference scale is given as
@@
The win estimands within other principal strata can be defined similarly. In the context of noncompliance, Mao (2024) defined a local Mann-Whitney treatment effect estimand given by $\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)|S_{1}=S_{2}=10\}$, which is also a special case of the PGCE. Different from our approach, Mao (2024) used an instrumental variable approach and discussed point and interval identification of this estimand based on monotonicity and exclusion restriction.